OD calibration fluctuations

Hi everyone,

Sometimes the Pio OD measurements start really high. For instance, a culture which was measured by spectrophotometer to have an OD=0.02, gets measured as OD=0.5 in the Pio, such as the case in the picture below:

What is wierd is that this behaviour disappears and comes back again. For instance, yesterday I started an experiment with two Pios which had a good start with an expected OD values as you can see below:

2 hours later, I unplugged and replugged the same Pios (due to change of location) to start a new experiment with a new culture which I measured with spectrophotometer to have an OD of 0.02. I did not change anything other than the location, so the config.ini, calibration, etc were the same. However, this time the same Pios:

  1. Started with a high OD measurement (~0.3)
  2. Showed use_calibration=0 measurement pattern

Hence, I have been thinking some potential causes, and how to troubleshoot the issue, such as:

  1. Could unplug - replug of workersre-states the config.ini file and accidentally sets use_config=0? If so would that be shown in the congi.ini? Because I personally checked and confirmed that use_config=1

  2. I have many Pioreactors on the same extension unit. Could this be caused due to power shortage?

  3. Could this be a software bug/error I managed to create in the most recent update, and if so does going back to 24.4.3 version (because we also encountered similar issue in version 24.4.11) can potentially solve the problem? Just to be clear, I am not trying to state this is a software update issue, this suggestion is based more on the personal observations.

  4. I was trying to change the voltage values on the OD calibrations of workers manually, perhaps this created problem on how to interpret the OD data. If so, could re-flashing the leader can potentially resolve the issue?

If anyone encountered a similar problem, any idea would be appreciated!